Problem #5

Data Table #2

Amount

$       1,000,000

Home Currency

Interest Home

5%

Interest Abroad

3%

Spot Rate

1.5500

FC/$

180-day Forward Rate

1.5000

FC/$

  1. By using data from Data Table #2 determine if the Market is in equilibrium by forecasting the future FX rate (FC/$) implied by the interest rates according to the interest rate parity.

Answer: 1.5349

  1. By using data from Data Table #2 compute Covered Interest Arbitrage profit or loss by investing in the Foreign Currency for 180 days. (Round your answer to the nearest dollar)

Answer: Profit of $23,833
Answer #1

Forward rate as per Interest rate parity = Spot rate(1+Interest rate FC)/(1+Interest rate $)

= 1.5500(1+3%*180/360)/(1+5%*180/360)

= FC 1.534878/$
i.e. FC 1.5349/$

Convert into FC today 1,000,000*1.55 = FC 1,550,000

Invest for 180 days and get 1,550,000(1+3%*180/360) = FC 1,573,250

Convert back into USD = 1,573,250/1.5 = $1,048,833.33

If invested in $, amount = 1,000,000*(1.025) = $1,025,000

Arbitrage Profit = $23,833.33